OpenStudy (anonymous):

The proportion below shows values for y to x that form a direct variation. What is the constant of variation? 6 divided by 2 equals twenty-four divided by 8 2 3 4 6

OpenStudy (anonymous):

Your on your own on this one sorry love♥

OpenStudy (anonymous):

this can be done by urslf only......

OpenStudy (anonymous):

can i help... did u get it

OpenStudy (anonymous):

3

OpenStudy (johnweldon1993):

Direct Variation has the formula \[\large y = kx\] To solve for the 'k' constant of variation...we divide both sides by 'x' \[\large k = \frac{y}{x}\] now we can plug in our numbers "6 divided by 2" \[\large k = \frac{6}{2} = 3\] "equals twenty-four divided by 8" \[\large k = \frac{24}{8} = 3\] The constant of variation = 3

OpenStudy (anonymous):

correct so am i

OpenStudy (anonymous):

guys i have been done with this basic math lmao